A solution is 0.126 M in KCl, 0.197 M in NaCl, and 0.148 M in MgCl2. What are the molarities of K + (aq), Na+ (aq), Mg2+ (aq), and Cl- (aq) in the solution? Hint: these are all ionic solutes and the salts fully dissociate into ions. Add up the concentrations of each ion that stem from each individual salt and then add them up as appropriate.
Answer #1

These are all ionic solutes and the salts fully dissociate into ions

KCl is ionize as follows:

KCl (aq)= K+(aq) + Cl-(aq)

0.126 M KCl will give 0.126 M   K+ and Cl-

NaCl (aq) = Na+(aq)+ Cl-(aq)

0.197 M NaCl will give 0.197 M Na+ and 0.197 M Cl-

MgCl2(aq) = Mg2+(aq)+ 2 Cl-(aq)

0.148 M MgCl2 will give 0.148 M Mg2+ and 2* 0.148= 0.296 M Cl-

Thus total Cl- =0.126 M   +0.197 M Cl-+0.296 M Cl-

= 0.619 M Cl-

The molarities of K + (aq), Na+ (aq), Mg2+ (aq), and Cl- (aq) in the solution are as follows:

0.126 M   K+

0.197 M Na+

0.148 M Mg2+

0.619 M Cl-